OpenStudy (anonymous):

Can you write a radical equation with an extraneous solution and show me how to solve it? or how to find that extraneous solution? Pleasee help!

OpenStudy (anonymous):

extraneous solutions are answers to an equation but not real roots.. and a real root is where the graph crosses the X axis

OpenStudy (tkhunny):

Just make one up. (x+1)(x-1) = 0 has solutions. \(\sqrt{x+1} = \dfrac{1}{\sqrt{x-1}}\) is a bit more difficult.